OpenStudy (anonymous):

(Tan^2x/(secx+1))+1???? ive already changed everything to dine and cosine and somehowgot cscx - 1 and it should be just 1 :/

OpenStudy (turingtest):

\[\sin^2x+\cos^2x=1\]divide this by \(\cos^2x\) and what do you get?

OpenStudy (anonymous):

Could you try and send that again please? It got corrupted...